### v2

# Taming Self-Training for Open-Vocabulary Object Detection

Abstract
Recent studies have shown promising performance in open-vocabulary object detection (OVD) by utilizing pseudo labels (PLs) from pretrained vision and language models (VLMs). However, teacher-student self-training, a powerful and widely used paradigm to leverage PLs, is rarely explored for OVD. This work identifies two challenges of using self-training in OVD: noisy PLs from VLMs and frequent distribution changes of PLs. To address these challenges, we propose SAS-Det that tames self-training for OVD from two key perspectives. First, we present a split-and-fusion (SAF) head that splits a standard detection into an open-branch and a closed-branch. This design can reduce noisy supervision from pseudo boxes. Moreover, the two branches learn complementary knowledge from different training data, significantly enhancing performance when fused together. Second, in our view, unlike in closed-set tasks, the PL distributions in OVD are solely determined by the teacher model. We introduce a periodic update strategy to decrease the number of updates to the teacher, thereby decreasing the frequency of changes in PL distributions, which stabilizes the training process. Extensive experiments demonstrate SAS-Det is both efficient and effective. SAS-Det outperforms recent models of the same scale by a clear margin and achieves 37.4 AP and 29.1 AP on novel categories of the COCO and LVIS benchmarks, respectively.
1 Introduction
Traditional closed-set object detectors [6, 16, 39] are restricted to detecting objects with a limited number of categories. Increasing the size of detection vocabularies usually requires heavy human labor to collect annotated data. With the recent advent of strong vision and language models (VLMs) [26, 36], open-vocabulary object detection (OVD) [14] provides an alternative direction to approach this challenge. Typically, OVD detectors are trained with annotations of base categories and expected to generalize to novel target categories by leveraging the power of pretrained VLMs.
One promising thread of recent studies for OVD [11, 12, 57, 49] leverages VLMs to obtain pseudo labels (PLs) beyond base categories. But they rarely explore self-training, a powerful and widely used schema for utilizing PLs in closed-set tasks [46, 45, 51, 52]. We investigate self-training for OVD and find the vanilla self-training approach does not improve OVD performance due to the following challenges.
First, the typical self-training in closed-set tasks sets a confidence threshold to remove noisy PLs based on the fact that the quality of PLs is positively correlated to their confidences. However, VLMs employed in OVD are pretrained for image-level alignment with texts instead of instance-level object detection that requires the localization ability. Thus, the confidence score from pretrained VLMs is usually not a good indicator for the quality of box locations (i.e., pseudo boxes) provided by PLs. For example, prior studies [14, 57] show that CLIP [36] tends to output imperfect object boxes as predictions with high confidence. Recent methods for OVD [57, 12] just apply thresholding to VLMs’ confidence scores and ignore the poor quality of pseudo boxes, which provides noisy supervision to the model. This issue becomes even worse when self-training is applied directly, since the noise accumulates which degrades the performance on novel categories. Moreover, these methods handle noisy PLs in the same way as ground truth of base categories during training, which further decreases the performance on base categories [12, 29].
Second, self-training for closed-set object detection [46, 52, 25] usually follows an online teacher-student manner. In each training iteration, the teacher generates PLs, and the student is trained with a mixture of ground truth and PLs. Then, the teacher is updated by the student with exponential moving average (EMA). However, we find such EMA updates degrade OVD models (see Table 4). Our hypothesis is that, unlike closed-set tasks, OVD provides no ground truth for target categories, and thus, the supervision for target categories is fully decided by the distribution of PLs predicted by the teacher. Hence, the EMA updates change the distribution of PLs in each iteration, making the training unstable.
In this paper, we propose Self-training And Split-and-fusion head for open-vocabulary Detection (SAS-Det) to tame self-training for OVD. First, we present a split-and-fusion (SAF) head to handle the noise of PLs. The SAF head splits the standard detection head into two branches: the closed-branch and the open-branch, which are fused at inference. The closed-branch, akin to the standard detection head, comprises a classification module and a box refinement module. It is supervised solely by ground truth from base categories, mitigating the impact of noisy PLs on the performance of base categories. The open-branch is a classification module supervised by class labels of both ground truth and PLs. It acquires complementary knowledge to the closed-branch and can significantly boost the performance when fused with the closed-branch. Moreover, this design circumvents noisy locations of pseudo boxes, reducing the accumulation of noise during self-training.
Second, instead of adopting the vanilla EMA update, we reduce the number of the updates and periodically update the teacher by the student. The quality of our PLs improves along with the periodic updates, and our final PLs are better than those of prior PL-based methods [12, 57] that introduce external handcrafted steps. Fig. 1 shows the key differences.
The proposed SAS-Det outperforms recent OVD models of the same scale by a clear margin on two popular benchmarks, i.e., COCO and LVIS. Without extra handcrafted steps, our pseudo labeling is more efficient than prior methods, i.e., nearly 4 times faster than PB-OVD [12] and 3 times faster than VL-PLM [57]. Extensive ablation studies demonstrate the effectiveness of the proposed components. On COCO, the thresholding of vanilla self-training decreases the performance of novel categories by 3.6 AP, and the EMA update decreases the performance by 6.9 AP. Instead, SAS-Det eliminates the degradation with two separate detection heads and the periodic update. The fusion of the SAF head boosts the performance by 6.0 AP.
The contributions of this work are summarized as follows. (1) We show two challenges of applying self-training to OVD and propose two simple but effective solutions, i.e., using different detection heads to mitigate the noise in PLs, and using periodic updates to reduce frequency of changes in PLs’ distributions. (2) The proposed SAF head for OVD handles the noisy boxes of PLs and enables fusion to improve the performance. (3) We present the leading performance on COCO and LVIS under widely used OVD settings and provide detailed analysis of the proposed SAS-Det.
2 Related Work
Vision-language models (VLMs). VLMs are trained to learn the alignment between images and text in a common embedding space. CLIP [36] and ALIGN [18] use contrastive losses to learn such alignment on large-scale noisy image-text pairs from the Internet. ALBEF [26] introduces multi-modal fusion and additional self-supervised objectives. SIMLA [21] employs a single stream architecture to achieve multi-level image-text alignment. FDT [8] learns shared discrete tokens as the embedding space. These VLMs achieve impressive zero-shot performance on image classification. But due to the gap between the pretraining and detection tasks, VLMs have limited abilities in object detection. In this work, we attempt to close the gap via PLs. There are studies [19, 27, 41] focusing on aligning any text phrases with objects. But they require visual grounding data that are more expensive than detection annotations. Our work scales up the vocabulary size for object detection without requiring such costly data.
Open-vocabulary object detection (OVD). Traditional closed-set object detectors [38, 39, 16, 6] achieve great performance but are limited in their vocabulary size. Zero-shot methods [3, 37, 60, 53] increase the vocabulary size but with limited accuracy. Motivated by the strong zero-shot abilities of VLMs, recent efforts focus on OVD. Finetuning-based methods [56, 33, 58, 24, 22] add detection heads onto pretrained VLMs and then finetune the detector with concepts of base categories. Such methods are simple but may forget the knowledge learned in the pretraining [2]. Distillation-based methods [14, 10, 48] introduce additional distillation loss functions that force the output of a detector to be close to that of a VLM and thus avoid forgetting. However, since the distillation losses are not designed for the detection task, they may conflict with detection objectives due to gradient conflicts that are a common issue for multi-task models. Methods like Gao et al. [12], Zhao et al. [57], Feng et al. [11], Wu et al. [49] create pseudo labels (PL) of novel concepts as supervision, and do not require extra losses, which sidesteps both catastrophic forgetting and gradient conflicts. But these methods need handcrafted steps to generate high quality PLs, e.g. multiple runs of box regression [57], activation maps [12] from Grad-CAM [42], image retrieval [11], or multi-stage training [49]. In this work, we address the two challenges of using self-training for OVD and enable an efficient end-to-end pseudo labeling pipeline. Besides, we point out the noise due to the poor locations of PLs and introduce the SAF head to handle such noise.
Self-training for object detection. Weakly-supervised object detection methods [46, 47, 40] explore online self-training by distilling the knowledge from the model itself. Recent semi-supervised object detection methods [52, 32, 25] adopt a teacher-student design, where the teacher is an exponential moving average (EMA) of the student. Self-training has been widely explored in the above fields but rarely in OVD. Unlike semi-supervised object detection, OVD encounters two challenges to use self-training, i.e., more noisy PLs, and larger changes in PLs’ distributions. This work proposes the SAF head to address the noise, and adopts periodic updates to reduce the frequency of changes in PLs’ distribution.
3 Approach
In open-vocabulary detection, an object detector is trained with bounding boxes and class labels of base categories . At inference, the detector is used for detecting objects of open concepts including and novel categories , where . To make such detection practical, recent studies [56, 14, 11, 29] adopt extra data (e.g., image-text pairs and image-level tags) and/or external VLMs. We follow their settings and leverage the pretrained CLIP [36] to build our OVD detector.
3.1 Adapting CLIP to OVD
In this section, we introduce how to adapt the ResNet-based CLIP into a Faster-RCNN [39] (C4) detector. For simplicity, we do not use FPN [31], but it can be incorporated with learnable gating from Flamingo [1]. Similarly, ViT-based CLIP models can be adapted to detectors following [28].
Region proposals from an external RPN. CLIP is pretrained via image-text alignment. As shown in Sect. 4.3, finetuning pretrained backbones to get region proposals decreases the performance, probably because such finetuning breaks the image-text alignment learned in the pretraining. To address this problem, F-VLM [24] freezes the pretrained backbone and only finetunes detection heads. But this solution limits the capacity of the detector. Unlike F-VLM, we follow the approach of Singh et al. [44], Zhong et al. [58] and employ an external RPN to generate region proposals, which is trained with ground truth boxes of base categories. Prior studies adopt RPN mainly to accelerate inference [44] or to improve region recognition [58]. By contrast, we aim to leverage the RPN to preserve the knowledge learned in the pretraining for better self-training.
Text embeddings as the classifier. For the -th region proposal from the external RPN, we apply RoIAlign [16] on the 4th feature maps of CLIP’s ResNet to get the proposal features. Then, the features are fed to the last ResNet block and the attention pooling of CLIP to get the region embedding , which is later used for classification. Following prior studies [14, 58, 29], we convert a set of given concepts with prompt engineering into CLIP text embeddings, which act as classifiers, as shown in Fig. 2b. A fixed all-zero embedding is adopted for the background category. Assuming is the embedding of the -th category, the probability of to be classified as the -th category is,
where denotes the cosine distance, denotes the vocabulary size, and denotes the temperature. In our detector, may be further fed to a box refinement module to predict the box shift based on the region proposal. With the above adaptions, our initial detector gains zero-shot detection ability to some extent. Please refer to the supplement for quantitative evaluations.
3.2 Taming Self-Training
Although self-training has been widely explored in closed-set object detection, using it for OVD presents two challenges, i.e., noisy PLs and frequent changes of PLs’ distributions. In the following, we describe our self-training pipeline and how to address the two challenges.
Self-training pipeline. Fig. 2a illustrates the pipeline of our self-training with a teacher-student manner. Training images are first fed into the RPN to obtain region proposals. Then, the teacher model runs inference on those proposals, where the resulting predictions with confidences above a threshold are selected as PLs. The student model adopts the same region proposals, and is supervised with both base ground truth and PLs generated by the teacher. The teacher is periodically updated with the parameters of the student model.
Handling noise in PLs. The noise of PLs from VLMs exists in both classification and localization. It is a common practice to filter PLs with classification confidence, but this only addresses noise in classification. To reduce the noise in locations of PLs, we exclude the noisy boxes of PLs from the training. That is, the classification loss is calculated on class labels of both ground truth and PLs, but the box regression loss is calculated only on ground truth boxes. Besides, to mitigate the impact of PLs on the performance of base categories, we propose SAF head and elaborate it in Sect. 3.3.
Reducing changes in PLs’ distributions with periodic updates. The teacher model is updated with the student to enable self-training. The exponential moving average (EMA) is a widely used approach for object detection [25, 32, 52], which updates the teacher in every iteration. But we observed empirically that it does not benefit OVD (Table 4). We hypothesize that, unlike semi-supervised object detection, OVD has no ground truth for the target categories. Thus, the distribution of the target data (i.e., PLs) is fully determined by the teacher. The EMA update changes the PLs’ distribution in each iteration, resulting in unstable training. As a solution, we periodically update the teacher after a set number of iterations to maintain consistent distributions for PLs between updates. We call this strategy as the periodic update and show it outperforms the EMA update by a large margin in Sect. 4.3.
3.3 Split-and-Fusion (SAF) Head
Our SAF head first splits a detection head into two branches, i.e. “closed-branch” and “open-branch”, with the goal to better handle noisy PLs during training. At inference, predictions from both heads are fused to boost the performance.
Splitting the detection head. The closed-branch follows a standard detection head with a classification module and a class-agnostic box refinement module. The former module classifies region proposals based on Eq. 1, and the latter one refines the proposal boxes for better locations. We train the closed-branch with boxes and class labels of ground truth for using standard detection losses, which include a cross entropy loss for classification and a box regression loss for localization. Since no PLs are used to train the closed-branch, the noise of PLs is unlike to impact its performance on . Moreover, as shown in Gu et al. [14], Zareian et al. [56], box regression modules trained on can generalize to novel categories that are unseen during training. Therefore, the closed-branch is able to provide generalized boxes, as well.
The open-branch only contains a classification module. It is trained using only the cross-entropy loss with class labels from both and PLs, hence, learning broader concepts beyond . Unlike distillation losses [14, 10], all losses for the two branches are originally designed for detection and are unlikely to conflict with each other. When generating PLs, we use the classification scores from the open-branch and bounding boxes from the closed-branch. Boxes of PLs are not directly used in our losses but are involved to select foreground proposal candidates for the classification loss.
Fusing complementary predictions. The open- and closed-branches are trained in different ways and learn complementary knowledge. Therefore, we fuse their predictions with the geometric mean at inference time. Specifically, assuming and are prediction scores of the open- and closed-branches, respectively, the final score is calculated as
where balances the two branches. The indices are the same as in Eq. 1. We keep a list of known base categories and take other categories beyond the list as novel. It’s important to note that fusion is a common strategy. For instance, F-VLM [24] employs score fusion to enhance the location quality of CLIP’s predictions. In contrast, our approach focuses on fusing the complementary predictions from the two branches. The primary contribution of our method lies in learning what to fuse rather than the fusion process itself.
4 Experiments
Method | Training Setup | Backbone | Detector | AP | AP | AP |
ViLD [14] | 16+LSJ | RN50-FPN | FasterRCNN | 27.6 | 59.5 | 51.2 |
VL-PLM [57] | 1+Default | RN50-FPN | FasterRCNN | 32.3 | 54.0 | 48.3 |
F-VLM [24] | 0.5+LSJ | RN50-FPN | FasterRCNN | 28.0 | - | 39.6 |
OV-DETR [55] | (Not Given) | RN50 | Deform. DETR | 29.4 | 61.0 | 52.7 |
CORA [49] | 3+Default | RN50 | DAB-DETR | 35.1 | 35.5 | 35.4 |
OVR-CNN [56] | (Not Given) | RN50-C4 | FasterRCNN | 22.8 | 46.0 | 39.9 |
RegionCLIP [58] | 1+Default | RN50-C4 | FasterRCNN | 26.8 | 54.8 | 47.5 |
Detic [59] | 1+Default | RN50-C4 | FasterRCNN | 27.8 | 51.1 | 45.0 |
PB-OVD [12] | 6+Default | RN50-C4 | FasterRCNN | 30.8 | 46.1 | 42.1 |
VLDet [29] | 1+LSJ | RN50-C4 | FasterRCNN | 32.0 | 50.6 | 45.8 |
BARON [48] | 1+Default | RN50-C4 | FasterRCNN | 33.1 | 54.8 | 49.1 |
SAS-Det (Ours) | 1+Default | RN50-C4 | FasterRCNN | 37.4 | 58.5 | 53.0 |
Method | Training Setup | Backbone | Detector | AP | AP | AP | AP |
ViLD [14] | 16+LSJ | RN50-FPN | FasterRCNN | 16.7 | 26.5 | 34.2 | 27.8 |
DetPro [10] | 2+Default | RN50-FPN | FasterRCNN | 20.8 | 27.8 | 32.4 | 28.4 |
F-VLM [24] | 9+LSJ | RN50-FPN | FasterRCNN | 18.6 | - | - | 24.2 |
BARON [48] | 2+Default | RN50-C4 | FasterRCNN | 17.3 | 25.6 | 31.0 | 26.3 |
SAS-Det (Ours) | 2+Default | RN50-C4 | FasterRCNN | 20.9 | 26.1 | 31.6 | 27.4 |
ViLD [14] | 16+LSJ | RN152-FPN | FasterRCNN | 19.8 | 27.1 | 34.5 | 28.7 |
OWL-ViT [33] | 12+LSJ | ViT-L/14 | OWL-ViT | 25.6 | - | - | 34.7 |
F-VLM [24] | 9+LSJ | RN50x4-FPN | FasterRCNN | 26.3 | - | - | 28.5 |
CORA [49] | 3+Default | RN50x4 | DAB-DETR | 22.2 | - | - | - |
SAS-Det (Ours) | 2+Default | RN50x4-C4 | FasterRCNN | 29.1 | 32.4 | 36.8 | 33.5 |
4.1 Experiment Setup
Datasets. We conduct experiments on the two popular OVD benchmarks based on COCO [30] and LVIS [15]. For the COCO dataset, 65 out of 80 categories are divided into 48 base categories and 17 novel categories. We report results on the validation set. This setting is denoted as COCO-OVD. For the LVIS dataset, 337 rare categories are used as the novel concepts and the remaining 866 categories (frequent and common) as the base concepts. We report results on the whole LVIS validation set. This setting is denoted as LVIS-OVD.
Evaluation metrics. Following prior studies [56, 14, 58, 48], we report AP at IoU threshold of 0.5 for COCO-OVD. AP, AP and AP are the metrics for the novel, the base and all (novel + base) categories, respectively. For LVIS-OVD, we report the mean AP averaged on IoUs from 0.5 to 0.95. AP, AP, AP and AP are the metrics for rare, common, frequent, and all categories.
Implementation details. We built our detector on Faster RCNN with the CLIP version of ResNet (RN50-C4 or RN50x4-C4) as the backbone. Following prior studies [55, 57, 59], we assumed the novel concepts were available during training and took them as our open concepts, but no annotations of novel concepts are used. Our method was implemented on Detectron2 [50] and trained with 8 NVIDIA A6000 GPUs. We trained our models with the 1 schedule (90k iterations) for COCO-OVD and the 2 schedule for LVIS-OVD. The batch size was 16 with an initial learning rate of 0.002. The default data augmentation of Detectron2 was applied. Loss terms were equally weighted. By default, the teacher models were updated three times, usually together with the decreases of the learning rate.
4.2 Comparison with the Existing Methods
COCO-OVD. We compare our method with prior work on COCO in Table 1. With the same backbone and detector, SAS-Det outperforms the most recent method BARON [48] by 4.3 AP on novel categories and 3.7 AP on base categories. Probably, BARON distills knowledge from VLMs, but distillation may lead to gradient conflicts and degrade the performance [57]. Compared to VLDet [29], SAS-Det gains an improvement of 7.9 AP. It is likely that VLDet employs noisy pseudo labels to train the detection head, impacting the detection of base classes in both classification and localization. In contrast, the two-branch design of our SAF head allows to train the closed-branch with just ground-truth of base classes and reduces the impact of the noise from pseudo labels. For more analysis on where our improvements come, please refer to Sect. 4.3 and Table 3. The first block of Table 1 provides results for methods with stronger backbones or detector architectures. When compared with those methods, SAS-Det still achieves the leading performance. Those results clearly demonstrate the effectiveness of SAS-Det.
LVIS-OVD. We provide the main results on LVIS in Table 2. When using ResNet50 as the backbone, SAS-Det achieves similar performance as the recent method DetPro [10]. DetPro proposes learnable prompts to generate better text embeddings as the classifier for OVD. Our method is orthogonal and can leverage DetPro’s prompts for further improvement. The first block of Table 2 also shows that SAS-Det outperforms ViLD by a large margin on novel categories (indicated by AP) but gets lower AP for base categories. This is also observed in the recent method BARON [48]. The performance gap is probably due to the training setup. ViLD is trained for 8 times more iterations and adopts a strong data augmentation method [13], both of which benefit detection on base categories. In contrast, BARON and our approach adopt shorter training and standard data augmentations. When replacing the pretrained CLIP ResNet50 with ResNet50x4 [36] as the backbone, we improve AP by 8.2 from 20.9 to 29.1, which demonstrates SAS-Det scales up nicely with stronger pretrained VLM backbones.
Ablation | AP | AP | AP |
Baseline | 31.4 | 55.7 | 49.4 |
(1) No external RPN, train the backbone for region proposals | (-6.0) 25.4 | 53.4 | 46.1 |
(2) Noisy boxes of PLs as supervision for box regression | (-3.6) 27.8 | 55.2 | 48.0 |
(3) No PLs, train with base data only | (-8.2) 23.2 | 56.9 | 48.1 |
(4) W/ SAF head, use closed-branch’s predictions ( in Eq. 2) | (-5.8) 25.6 | 57.9 | 49.5 |
(5) W/ SAF head, use open-branch’s predictions ( in Eq. 2) | (+0.5) 31.9 | 57.4 | 50.7 |
(6) W/ SAF head, use fused predictions ( in Eq. 2) | (+6.0) 37.4 | 58.5 | 53.0 |
Update strategy | AP | AP | AP |
Baseline (w/ periodic update) | 31.4 | 55.7 | 49.4 |
(7) No teacher and no pseudo labels | (-8.2) 23.2 | 56.9 | 48.1 |
(8) Not update the teacher | (-1.8) 29.6 | 55.9 | 49.1 |
(9) Take the EMA of the student as the teacher | (-6.9) 24.5 | 53.9 | 46.2 |
(10) Replace the teacher with the student every iteration | (-23.4) 8.0 | 53.3 | 41.5 |
# Update | AP |
8 | 27.6 |
4 | 30.6 |
3 (baseline) | 31.4 |
2 | 31.6 |
1 | 30.9 |
0 (No update) | 29.6 |
Method | AP | AP | AP |
OVR-CNN [56] | 11.5 | 22.9 | 38.1 |
Detic-80 [59] | 11.5 | 27.3 | 38.3 |
MEDet [7] | 18.6 | 32.6 | 42.4 |
Ours | 23.0 | 37.1 | 47.2 |
4.3 Ablation Studies
In this section, the default baseline shares the same architecture and the training as our SAS-Det (with RN50-C4 as the backbone), except that the SAF head is replaced with a single detection head. The localization module of the detection head is trained with base ground truth boxes only, but the classification module is trained with class labels of both ground truth and PLs. Thus, the baseline provides a naive solution to exclude noisy pseudo boxes from the training. All evaluations are conducted on COCO.
External RPN. We leverage an external RPN to generate region proposals so that finetuning focuses on region-text alignment that is similar to the pretraining task. In this way, the detector is unlikely to forget the knowledge obtained in the pretraining. To validate the effectiveness of the external RPN, we follow F-VLM [24] to train a detector without an external RPN (See the supplement for more details). Table 3 compares baseline with the detector in the row of (1). As shown, without the external RPN, the performance drops on both novel and base categories.
Removing boxes of PLs from training. Our baseline handles location noise of PLs’ boxes by directly excluding them from the training. As shown in Table 3, baseline outperforms (2)’s detector. The only difference between them is the removal of pseudo boxes from training. This result clearly shows that simply removing the pseudo boxes is an effective way to deal with location noise.
Splitting the detection head. As shown in Table 3, compared to baseline, the open-branch of the proposed SAF head in (5) achieves similar performance on novel categories. This indicates that the split handles the noise in PLs’ location as well as the naive solution of baseline. This is plausible because both solutions exclude the noisy boxes of PLs from training.
Additionally, the open-branch gains 1.7 AP on base categories. Compared to (3)’s, the closed-branch in (4) gain improvements in terms of both AP and AP. Note that (4)’s and (3)’s heads are trained with the same data. Based on those results, the split benefits OVD beyond handling the location noise. One advantage is that, due to the split, the closed-branch is less likely to be influenced by the noise of PLs and learn how to better localize objects.
Fusing predictions. Our SAF head fuses predictions of the open-branch and the closed-branch at inference time. As shown in Table 3, (6)’s outperforms all the others by a large margin with the prediction fusion. Note that (4)’s, (5)’s and (6)’s refer to the different outputs of the same model. The improvement can be attributed to the fact that the two branches are trained on different sets of data and learn complementary knowledge. The closed-branch is trained with class labels and boxes of base categories. The open-branch is trained with class labels of more vocabularies including base and novel classes. Thus, the former learns more about how to localize objects and how to detect base classes. The latter learns more about how to detect new objects, which complements the former.
Update strategies for the teacher model. The teacher is updated with the student model to improve PLs during finetuning. We evaluated several strategies for updating the teacher model in Table 4 and summarize our findings as follows. First, the EMA update in (9), which is shown effective and widely used in semi-supervised object detection [52, 32, 25], is just slightly better than training without PLs in (7). It is worse than no update to the teacher (i.e. no self-training) in (8). Second, if we replace the teacher with the student every iteration as (10), AP drops to 8.0 that is much lower than that number in (7). These findings indicate that it’s harmful to change the teacher frequently. We hypothesize that frequent updates change the distribution of PLs too often and makes the training unstable. Last but not least, by reducing the number of updates, our periodic update significantly outperforms the EMA update in OVD.
The number of updates to the teacher model. We trained our detectors with different numbers of updates to teacher models. Generally, we consider the learning rate schedule and distribute updates as even as possible during training. Please see the supplement for more details about how we distribute the updates. As shown in Table 6, too many updates, e.g., 8 or 4 updates, lead to performance drops mainly due to the following. First, similar as the aforementioned EMA update, too many updates change the distribution of PLs too often and make the training unstable. Second, the more updates, the earlier an update happens. However, the student model is not well trained at the early stage of the training and thus is not good enough to update the teacher. Table 6 shows that 2 and 3 updates achieve similar performance. But we set 3 updates as default to include as many updates as possible, considering that the only overhead is to update the weights of the teacher with the student.
4.4 Further Analysis
Evaluation on the retained 15 COCO categories. Following prior studies [55, 57, 59], SAS-Det assumes novel concepts are available during training and achieves good performance. A natural question is if SAS-Det generalizes to alien concepts that are unknown before the evaluation. To answer the question, we follow the evaluation protocol of MEDet [7] where an OVD detector is evaluated on the whole COCO validation set with all 80 COCO concepts. AP, AP and AP are reported as APs averaged on 15 retrained, 17 novel and all 80 COCO categories, respectively. As shown in Table 6, SAS-Det achieves leading performance by a clear margin, which indicates the good generalization capability of the proposed method.
PL’s quality during self-training. We compare our PLs and the prior state-of-the-art PLs from VL-PLM [57] on the COCO validation set in terms of PLs’ quality. Following VL-PLM, we take AP as the metric and evaluate our teacher models during training. As illustrated in Fig. 3, the quality of our initial PLs is not as good as VL-PLM’s. But ours get close to VL-PLM’s after two updates and become the better by the third update. This is because the teacher is initialized with CLIP, which is not pretrained for detection, and thus cannot provide high quality PLs. With updates, the teacher is equipped with the knowledge about detection and generates promising PLs.
Time cost of pseudo labeling. We run two major PL-based methods [12, 57] on our computational environment and compare them with our pseudo labeling in terms of the time cost in Table 3. As shown, our method is 3 faster than VL-PLM [57] and nearly 4 faster than PB-OVD [12], offering better PLs.
Reducing the time cost of external RPN. There is a simple solution to remove the extra computational cost of external RPN adopted by SAS-Det. We first generate pseudo labels with SAS-Det, and then follow VL-PLM [57] to train a Faster RCNN based on ResNet50 for OVD without the external RPN (“Faster RCNN + PLs”). As shown in Table 8, “Faster RCNN + PLs” achieves slightly better or close performance as SAS-Det.
Method | AP | AP | AP |
SAS-Det | 37.4 | 58.5 | 53.0 |
Faster RCNN + PLs | 38.1 | 58.3 | 53.0 |
5 Conclusion
In this paper, we highlight two challenges associated with applying self-training to OVD: 1) noisy PLs from pretrained VLMs and 2) frequent changes of PLs’ distributions. To address the challenges, we introduce a split-and-fusion (SAF) head and implement periodic updates. The SAF head splits a standard detection head into an open-branch and a closed-branch. The open-branch is trained with the class labels of ground truth and PLs. The closed-branch is trained with both class labels and boxes of ground truth. This approach effectively eliminates the location noise of PLs during training. Furthermore, the SAF head acquires complementary knowledge from different sets of training data, enabling fusion to enhance performance. The periodic updates reduce the number of updates to the teacher models, thereby decreasing the frequency of changes in PLs’ distributions. We demonstrate that the proposed method SAS-Det is both efficient and effective. Our pseudo labeling is much faster than prior PL-based methods [12, 57]. SAS-Det also outperforms recent models of the same scale on both COCO and LVIS for OVD.
Supplementary Material
This document is supplementary to the main paper as,
Sect. A.1 elaborates how to train a detector for open-vocabulary object detection (OVD) without an external RPN, which is used in “External RPN” of Sect. 4.3 in the main paper.
Sect. A.2 shows the quality of pseudo labels (PLs) from the initial teacher and how to improve initial PLs for Sect. 3.1.
Sect. A.3 elaborates the exact iterations when we update the teacher model for “The number of updates to the teacher model.” of Sect. 4.3.
Sect. A.4 shows the performance of our method with different initial weights.
Sect. B.1 compares the proposed SAS-Det and visual grounding methods.
Sect. B.2 shows how well SAS-Det preserves knowledge learned in the pretraining.
Sect. B.3 provides extra ablation studies on LVIS.
Sect. B.4 explores stronger external RPNs for SAS-Det.
Sect. B.5 explores to use external PLs at the beginning of self-training
Sect. C.1 visualizes good and failure cases of PLs.
Sect. C.2 visualizes good and failure cases of our OVD detector.
Sect. D discusses some limitations of this work and potential solutions.
Appendix A Extra Details for The Main Paper
A.1 Our detector for OVD without an external RPN
This section explains how we train a detector without an external RPN. This approach is compared with the baseline in paragraph “External RPN” of Sect. 4.3. As shown in Fig. 4, we divide the training into two stages: (a) In the first stage, we put a RPN box head on top of a pretrained but frozen CLIP image encoder and only train the box head with a box loss. The RPN box head outputs region proposals with objectness scores. The box loss consists of a foreground classification term and a box regression term. (b) In the second stage, the RPN box head and a detection head are built on the same CLIP image encoder. No modules are frozen. The whole model is trained the same way as baseline in the main paper with online pseudo labels from a teacher model.
Table 10 provides the performance of the first and the second training stages. We evaluate the model of the 1st stage by directly classifying region proposals with text embeddings. The model of the 1st stage can be regarded as the initial teacher of the 2nd stage. It achieves similar performance as the initial teacher of baseline, which indicates that initial PLs of baseline and the 2nd stage share similar qualities. However, baseline outperforms the 2nd stage’s model on both base and novel categories. Such results clearly demonstrate that an external RPN is important to our detector and training.
Models | AP | AP | AP |
First stage (no external RPN) | 10.5 | 12.7 | 12.1 |
Second stage (no external RPN) | 25.4 | 53.4 | 46.1 |
Initial teacher of baseline (w/ an external RPN) | 11.9 | - | - |
baseline (w/ an external RPN) | 31.4 | 55.7 | 49.4 |
Method | AP (COCO) | AP (LVIS) |
w/o RPN scores | 3.8 | 1.9 |
w/ RPN scores | 11.9 | 9.1 |
A.2 Improving initial pseudo labels with RPN scores
In the self-training process, we initialize the teacher model with pretrained CLIP weights to generate PLs. However, the initial teacher cannot provide high quality PLs because CLIP is weak at localizing objects and has poor zero-shot detection ability [14, 57]. Similar to VL-PLM [57], we average the objectness scores from the external RPN with the prediction scores from the teacher model. Assuming is the objectness score of the -th region proposal, the averaged prediction score is where refers to the -th category. Table 10 provides the quantitative results for whether or not to use the RPN fusion. As shown, without the fusion, the quality of PLs significantly drops.
A.3 When to update the teacher model.
In this section, we discuss the timing of updates to the teacher model during training. We trained our detectors with different times of updates to teacher models and provides the exact iterations when we update the teacher model for COCO-OVD in Table 11. Generally, we consider the learning rate schedule and distribute updates as evenly as possible during training. As shown in Table 11, too many updates, e.g., 8 or 4 updates, lead to performance drops mainly due to the following. First, similar as the aforementioned EMA update, too many updates change the distribution of PLs too often and make the training unstable. Second, the more updates, the earlier an update happens. However, the student model is not well trained at the early stage of the training and thus is not good enough to update the teacher. Table 11 shows that 2 and 3 updates achieve similar performance. But we set 3 updates as default to include as many updates as possible, considering that the only overhead of our update is to copy the weights from the student to the teacher. For LVIS-OVD, we find that a later update helps and only conduct the update when the learning rate changes. For the 2 training setup, the teacher is updated at 120k and 160k iterations.
# Updates | Iterations to update | AP |
8 | Every 1k iterations | 27.6 |
4 | Every 2k iterations | 30.6 |
3 (baseline) | 4k,6k,8k | 31.4 |
2 | 4k,8k | 31.6 |
1 | 5k | 30.9 |
0 (No update) | N/A | 29.6 |
A.4 Initial weights.
We use CLIP weights from RegionCLIP [58] to initialize SAS-Det. However, we noticed that RegionCLIP’s pretraining includes LVIS base boxes that may include boxes for novel categories of COCO. To avoid potential data leakage, for our experiments on COCO-OVD, we followed RegionCLIP’s procedure to finetune CLIP with COCO base boxes only. Besides, On COCO-OVD, we report results with Soft NMS [4] as RegionCLIP [58] did.
Table 12 compares original CLIP weights with RegionCLIP weights as the initialization of baseline. As shown, initialized with either of them, the detectors achieve similar performance on novel categories on both the COCO and LVIS datasets. This is probably because our finetuning with PLs closes the gap between CLIP’s and RegionCLIP’s pretraining. Since RegionCLIP does not benefit our method, it is still fair to compare SAS-Det with other methods that uses the original CLIP. Table 12 also shows that initialization with RegionCLIP improves the performance on base categories. Since RegionCLIP adopts the boxes of base categories in its pretraining, it actually provides longer training on base categories. We attribute the improvement on base categories to the longer training.
Initial weights | COCO-OVD | LVIS-OVD | |||||
AP | AP | AP | AP | AP | AP | AP | |
From original CLIP | 31.1 | 53.5 | 47.7 | 19.6 | 23.6 | 30.6 | 25.6 |
From RegionCLIP pretraining | 31.4 | 55.7 | 49.4 | 19.8 | 25.3 | 31.5 | 26.8 |
Method | Training Data | Backbone | Detector | AP | AP | AP | AP |
MDETR | LVIS, GoldG+ | RN101 | DETR [6] | 7.4 | 22.7 | 25.0 | 22.5 |
GLIP | O365, GoldG, Cap4M | Swin-T | DyHead [9] | 20.8 | 21.4 | 31.0 | 26.0 |
X-DETR | LVIS, GoldG+, CC, LocNar | RN101 | Def.DETR [61] | 24.7 | 34.6 | 35.1 | 34.0 |
Ours | LVIS Base | RN50-C4 | FasterRCNN | 20.9 | 26.1 | 31.6 | 27.4 |
Ours | LVIS Base | RN50x4-C4 | FasterRCNN | 27.3 | 30.7 | 35.0 | 31.8 |
Appendix B Additional Experiments
B.1 Comparison with visual grounding models
There are some recent studies [20, 27, 5] focusing on large-scale pretraining for visual grounding where text phrases in a whole caption are aligned with objects. They usually train their models with multiple datasets, including detection data, visual grounding data [27], and image-text pairs. Visual grounding data requires the association between each box and each specific text phrase, which is much more expensive than detection annotations. The datasets GoldG+ and GoldG, introduced in [27],are two widely used Visual grounding datasets. GoldG+ contains more than 0.8M human-annotated gold grounding data curated by MDETR [20], including Flickr30K [34], VG Caption [23], GQA [17], and RefCOCO [54]. GoldG removes RefCOCO from GoldG+. Image-text pairs are usually from CC [43] and LocNar [35].
Table 13 compares our method with MDETR [20], GLIP [27] and X-DETR [5] on LVIS minival that contains 5k images. As shown, our detector with RN50x4 as the backbone achieves the leading performance on rare categories without using any annotations of those categories. By contrast, the other three methods employ the costly visual grounding data. MDETR [20] and X-DETR [5] even adopt the ground truth of rare categories but cannot outperform our method. Such results further demonstrate the effectivenss of the proposed SAS-Det.
B.2 Preserving the knowledge from the pretraining
In this section, we explore if models after our finetuning generalizes as well as pretrained models. If so, our finetuning preserves the knowledge learned in the pretraining. Specifically, we evaluated the proposed SAS-Det and baseline, which are trained on COCO-OVD with 65 concepts, on LVIS with 1203 concepts. Then, we compare them with the pretrained models and report the results in Table 14. As shown, though finetuned with limited concepts, SAS-Det and baseline achieve similar or better performance as pretrained models on rare categories of LVIS. Note that those categories do not appear during finetuning. This indicates that the knowledge learned in the pretraining is successfully preserved in our finetuning. We attribute the improvement on common and frequent categories to the fact that finetuning adapts CLIP to OVD and thus the models learn how to better handle instance-level detection instead of image-level classification that CLIP is pretrained for.
Method | AP | AP | AP | AP |
Original CLIP | 8.7 | 6.7 | 4.0 | 6.0 |
RegionCLIP pretraining | 9.7 | 7.1 | 4.3 | 6.4 |
baseline | 9.2 | 9.7 | 9.0 | 9.4 |
SAS-Det (Ours) | 13.1 | 10.5 | 9.9 | 10.7 |
B.3 Ablation studies on LVIS
In this section, we provide ablation studies on LVIS, which are similar as what we did on COCO. Except trained on LVIS-OVD, the baseline model LVIS baseline is the same as baseline in the main paper. As shown in Table 15, we have consistent observations on LVIS as on COCO. First, based on the results of LVIS baseline and (1)’s, it is beneficial to remove noisy pseudo boxes from finetuning. Second, (3)’s outperforms both LVIS baseline and (2)’s, which demonstrates that the proposed SAF head helps. The improvement probably comes from the fact that the SAF head avoids noisy pseudo boxes as supervision and incorporates a fusion from different branches.
Ablation | AP | AP | AP | AP |
LVIS baseline | 19.8 | 25.3 | 31.5 | 26.8 |
(1) Use noisy pseudo boxes to train box regression | (-4.5) 15.3 | 24.0 | 31.1 | 25.2 |
(2) No pseudo labels, train with base data only | (-3.1) 16.7 | 27.0 | 33.0 | 27.6 |
(3) SAF head, fuse the open- and the closed-branches | (+1.1) 20.9 | 26.1 | 31.6 | 27.4 |
Training boxes for RPNs | AP | AP | AP |
(4) COCO Base (48 categories) | 31.4 | 55.7 | 49.4 |
(5) COCO Base + Novel (65 categories) | 32.7 | 55.7 | 49.7 |
(6) COCO (80 categories) | 32.9 | 55.7 | 49.8 |
(7) LVIS Base (866 categories) | 32.9 | 55.2 | 49.3 |
Method | AP | AP | AP |
baseline | 31.4 | 55.7 | 49.4 |
baseline + VL-PLM’s PLs | 33.5 | 55.9 | 50.1 |
B.4 Inference with different RPNs
The proposed SAS-Det leverages an external RPN to get region proposals, and thus it is open to other RPNs without any further finetuning. As shown in Table 17, our model is evaluated together with several RPNs that are trained with different data. Based on those results, we have several findings. First, the model is only trained with (4)’s RPN but achieve similar or better performance with other RPNs. This indicates that our model does not rely on the specific RPN that is used in the training, and it is open to different RPNs. Second, (4)’s performance is close to others on novel categories, but its RPN is trained without any boxes of novel categories. This demonstrates that the RPN trained on base categories generalizes to novel ones.
B.5 Using external PLs at the beginning
At the beginning of self-training, our PLs are not as good as VL-PLM’s [57], but our training pipeline allows us to leverage external high quality PLs before the first update to the teacher. To demonstrate this, we trained baseline on COCO-OVD using VL-PLM’s PLs before the first update. As shown in Table 17, with VL-PLM’s PLs, our model improves from 31.4 to 33.5 for novel categories.
Appendix C Qualitative Results
C.1 Visualizations of pseudo labels
Figure 5 provides failure cases of our final PLs (after all updates) on the COCO dataset. We find two major types of failures. (a) Redundant boxes. In this case, one object has multiple predictions that are overlapped with each other. Those overlapped PLs indicate that the pseudo boxes are extremely noisy and cannot be improved by a simple thresholding based on classification confidences. Thus, it is necessary to handle the noise in pseudo boxes separately. We believe that redundant PLs are caused by the poor localization ability of CLIP that provides noisy initial PLs. Though our finetuning improve the localization ability to some extent, how to further improve this ability is still challenging and requires future research. (b) Wrong categories. We find the teacher model tends to classify every object into given concepts and generates PLs with wrong categories. Fortunately, there are usually some connections or similarities between the detected objects and the wrong categories. OVD employs text embeddings as classifiers, and text embeddings of related concepts share similarities. For example, the embeddings of “bus” and “train” are close to each other. Thus, though with wrong categories, those PLs may still provide supervision for OVD to some extend.
Figure 6 visualizes more PLs with different times of updates to the teacher model. All samples come from the COCO dataset. As shown, PLs before the update are noisy. Updates remove the noise and improve the quality of PLs.
C.2 Visualizations of our detector for OVD
We visualize good cases and failure cases of the final detector in Fig. 7 and Fig. 8, respectively. All samples come from the COCO dataset, and only predictions of novel categories are provided. As shown in Fig. 7, our detector is able to detect rare objects, e.g. a toy umbrella, a bus with rich textures, and an elephant sculpture.
We find two major failure cases as shown in Fig. 8. (a) Missing instances that usually happen in images with a crowd of objects belonging to one category. In our view, such cases are difficult for fully supervised object detection, let alone OVD. (b) Redundant predictions. We believe this is caused by using redundant PLs (see examples in Fig. 5a) as supervision. Improving PLs will alleviate redundancy in predictions.
Appendix D Limitations
This work has the following limitations that can be further investigated: (1) Compared to standard training, our self-training involves an additional teacher model, which requires more GPU memory. One possible solution to reduce the cost is to alternatively run the teacher and the student. (2) Although much faster than prior methods, our online pseudo labeling still induces overhead during training when millions of iterations are required. One possible solution is to generate offline PLs each time the teacher model is updated. (3) Although achieving good performance, SAS-Det still suffers from two major failure cases, which are visualized in Sect. C.2. The future work may explore stronger pretrained VLMs and better denoising steps for solutions.